OpenStudy (kanwal32):

how to prove n sigma r=1=n(n+1)/2

OpenStudy (kanwal32):

\[\sum_{r=1}^{n}r=n(n+1)/2\]

OpenStudy (kanwal32):

how to prove this

ganeshie8 (ganeshie8):

you may find this interesting :) http://mathcentral.uregina.ca/QQ/database/QQ.02.06/jo1.html

OpenStudy (kanwal32):

but how to prove it

OpenStudy (anonymous):

First use: \[ \sum r = 1+2+\dots +n = n + (n-1) + \dots +1 = \sum (n+1)-r \]

OpenStudy (anonymous):

Therefore \[ 2\sum r = \sum r + \sum(n+1)-r = \sum r + (n+1) - r = \sum n+1 = n(n+1) \]So \[ 2\sum r = n(n+1) \implies \sum r = \frac{n(n+1)}{2} \]